29 January 2008

RHIC pp Set-Up:  Maintenance/Set-Up for pp started at 8:00AM yesterday.  Power supply work went well, completing it in one "day".  Beam was brought to RHIC after midnight.  There was circulating beam in Blue and Yellow with good lifetime. 

Today there will be work on lifetime, RF capture, and Instrumentation. 

Beam out of the AGS has been 1.7E11, with a polarization at 1.5E11 of 40%.

STAR (pp2pp) has requested a 2 hour access during the Power Supply ramp development work for Don some time today.  Don's work will take ~4 hours. 

We need to bring on the Snakes in RHIC before RF capture.  Snake work will be scheduled first, from 9:00-12:00.  RF Capture will follow from 12:00-16:00.  Power Supply ramp development will follow 16:00-20:00. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020